HETAV PATEL

Grade 11,

What will be the answer to this question If Y=2at and if x=at^2Then d2y/dx2 is =

What will be the answer to this question If Y=2at and if x=at^2Then d2y/dx2 is =

Grade:9

1 Answers

sonika p
6 years ago
d2y/dx2 = (d2y/dt2)/(d2x/dt2)dy/dt = 2ad2y/dt2 =0dx/dt = 2atd2x/dt2 = 2aTherefore, d2y/dx2 = 0/2a = 0